Circular No. 6883

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

     URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 H1

On Apr. 22, the Central Bureau received a report from Patrick L. Stonehouse of an apparent visual comet discovery, with motion stated to the north (later to the northwest), and the object was described as diffuse with slight condensation. Poor initial positions evidently caused attempts at confirming the comet over several square degrees with CCDs by T. Kojima (Chiyoda) and M. Tichý (Kleť) on Apr. 22–23 UT, and by H. Mikuž and J. Skvarc (Crni Vrh) on Apr. 25, to be unsuccessful. Continuing observations by Stonehouse encouraged the Bureau to issue refined search ephemerides, resulting in confirmation as indicated below (observations from Japanese observers were communicated by S. Nakano, Sumoto):

1998 UT  α_{2000}  δ_{2000}  m_{1} Observer
Apr. 22.25  15^{h}54.^{m}6  + 0^{o}21′  12  Stonehouse
23.016  15 51.2  + 2 15  13  _{”}
24.229  15 44.0  + 5 20  _{”}
25.0  15 36.4  + 8 03  _{”}
26.265  15 30.1  +10 44  10.7 Hale
26.53337 15 28 43.45 +11 30 16.3  Garradd
26.53461 15 28 42.94 +11 30 28.0  _{”}
26.53586 15 28 42.44 +11 30 39.7  _{”}
26.53756 15 28 41.76 +11 30 55.8  _{”}
26.53865 15 28 41.30 +11 31 06.0  _{”}
26.61007 15 28 12.87 +11 42 01.5 11.5 Nakamura
26.61215 15 28 12.04 +11 42 21.4  _{”}
26.61424 15 28 11.19 +11 42 41.0  _{”}
26.67567 15 27 46.45 +11 52 18.0 12.4 Ikari
26.67660 15 27 45.79 +11 52 43.1  Garradd
26.67763 15 27 45.42 +11 52 53.4  _{”}
26.68012 15 27 44.64 +11 53 00.2  Ikari

P. L. Stonehouse (Wolverine, MI). 0.44-m reflector. A. Hale (Cloudcroft, NM). Visual observation with 0.41-m reflector. Coma

diameter 4′.5; comet diffuse with condensation; motion northwest. G. J. Garradd (Loomberah). 0.25-m f/4.1 reflector + CCD. Tail 30′′ long

in p.a. 155^{o} on short CCD exposures; m_{2} = 14.8–15.4. A. Nakamura (Kuma). Communicated by S. Nakano. 0.60-m f/6 reflector

+ CCD. Tail > 5′ long in p.a. 202^{o}, extends out of CCD field. Y. Ikari (Moriyama). 0.25-m f/6.3 reflector + CCD.

1998 April 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green